A 11.1 g sample of an aqueous solution of perchloric acid contains an unknown amount of the acid. If 29.9 mL of 0.534 M sodium hydroxide are required to neutralize the perchloric acid, what is the percent by mass of perchloric acid in the mixture?
1
Expert's answer
2019-10-21T07:16:39-0400

"HClO_4 + NaOH \u2192 NaClO_4 + H_2O"

By using Molarity equation,

"M_1V_1=M_2V_2" (1 for Acid,2 for NaOH)

Let "x" be mole fraction of "HClO_4" in the solution.

Molar mass of perchloric acid"=100.5\ gm"

So,"\frac{11.1}{100.5}\times x=29.9\times 10^{-3}\times 0.534"

"x=0.144563"

Mass percentage"=x\times 100=14.46" %